1.k2[Fe(Co)8]+CH2I2=?+KI
2.Mo2(ch3coo)4+hcl=?+ch3cooh+H(PLUS)
3.Mo(Co)6+CH3COOH=?+CO2+H2
4.CO+H2+CO=?
1
Expert's answer
2013-01-15T09:12:04-0500
1. K2[Fe(CO)8] + CH2I2 = CH2[Fe(CO)8] + 2KI

2. Mo2(CH3COO)4 + 8HCl = Mo2Cl84- + 4CH3COOH + 4H+ (temperature is 0 C)

3. Mo(CO)6 + 2CH3COOH = Mo(CH3COO)2 + 6CO + H2

4. CO + 2H2 = CH3OH